User Tools

Site Tools


lbaops:telecon2009feb11

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

lbaops:telecon2009feb11 [2009/02/12 15:12]
guest
lbaops:telecon2009feb11 [2015/12/18 16:39]
Line 1: Line 1:
-====== ​ VLBI teleconference,​ 11 February 2009 ====== 
-=====  12.00 Perth, 14.00 Eastern, 16.00 New Zealand ===== 
- 
-Phone: 
- 
-| Vic:|03 9545 8222 | 
-| NSW:|02 9490 8022 | 
-| ACT:|02 6276 6622 | 
-| Qld:|07 3214 2311 | 
-|  WA:|08 6436 8844 | 
-|  SA:|08 8273 8122 | 
-|  NT:|08 8944 8488 | 
-| Tas:|03 6232 5188 | 
-Meeting ID: 5007 
- 
-Chair: Tasso Tzioumis 
- 
-Attendance: 
-Tingay, Reynolds, Moin, Bignall, C. Hotan, A. Hotan (Curtin); 
-Tzioumis, Phillips (ATNF Marsfield); 
-Horiuchi (Tidbinbilla);​ 
-Lovell (Hobart); Reid (Mt. Pleasant); 
-Stevens (ATNF Narrabri); 
-Rioja, Dodson (UWA) 
- 
-**Agenda** 
- 
-  *  Review of [[lbaops:​telecon2009jan28 | Actions from last meeting]] 
- 
-  *  Next VLBI session ​ 
-    *  Schedules 
-    *  Disks (problems with 750G disks) 
-    *  e-VLBI tests 
-    *  Observing support 
- 
-  *  e-VLBI tests to Curtin ​ (<1 Gbps and >1 Gbps) (Phillips/​Reynolds) 
-  *  e-VLBI demo to Curtin (Tzioumis/​Tingay) 
-  *  LBA operations/​eVLBI/​Networking meeting (Tzioumis/​Bignall) 
-  *  Update on ASKAP VLBI EoI and ICRAR investment plan (Tingay) 
-  *  Status of ARCS data transfer tests to iVEC (Phillips) 
-  *  Update on New Zealand antenna (Gulyaev) 
-  *  Update on Curtin cluster at Narrabri (Phillips/​Amy) 
-  *  AuScope Update (Lovell) 
-  *  Update on VLBI tests with GMRT (Tzioumis) 
-  *  AOB 
-    *  Mopra recording problems from November (Phillips) 
- 
-**Actions arising** 
- 
-(new items in **blue**) 
- 
-GENERAL 
-  *  One of Curtin Mark5B+ units to be shipped to AUT when they arrive, plus an invoice to AUT. (Steven) - updated 28/01/09 - Shipping cockup by supplier. ​ Delay in delivery of units. 11/2/09 - in progress. 
- 
-  *  Swinburne archive of correlated VLBI data to be moved to ATNF & Curtin at end of all Swinburne correlation (Steven/​Claire) - update 28/01/09 - Swinburne correlation complete. ​ Claire transferring data to Curtin. ​ Will tidy up before sending on to ATNF.  Claire to advise ATNF on archive size. 11/2/09 - in progress. 
- 
-  *  Options for disks and data transport: 
-    *  Buying excess disks from Swinburne VLBI pool and/or Mk5 disks. (Tasso/​Steven) - update 28/01/09 - Tasso and Steven still discussing. 
-    *  Initiate discussion on next generation VLBI systems and disks (Mk5b/​Mk5c/​of-the-self) for the March LBA meeting. (Tasso/​Steven) - update 28/01/09 - to be added to agenda for March meeting. ​ UTas to be debriefed and kept in loop with meeting outcomes. 
- 
-  *  Provide additional advice to PIs on scheduling the ATCA to keep it phased up at 8 GHz and above (Tasso) - update 28/01/09 - Richard to have input regards polarisation properties. 11/2/09 - to be tested with observations in next session. 
- 
-  *  List on the Wiki available bands/​polarisations at each LBA antenna. Also update "​standard bands" (from S2) and define "​standard"​ bands for wide-band recording. (Cormac/​Tasso). 11/2/09 - initial progress within SCHED. Will be further followed up. 
- 
-  *  Curtin Cluster at Narrabri: we would like to get the cluster up once the network is installed. Action on Tasso and Shaun to get the networking going as soon as possible (Tasso/​Shaun) - update 28/01/09 - Gear ordered and arrived. ​ Work to be done in next ~3 weeks. ​ Jamie at Narrabri maybe get involved. 
- 
- 
-DEMO and MEETINGS in March 
- 
-  *  Tasso to prepare a 'block schedule'​ for the meeting and seek comment (Tasso) 
- 
-  *  **eVLBI tests/demos to be discussed at Mar 2 meeting with AARNet. ​ Agenda of issues needed for that meeting (Steven/​Tasso/​Chris).** 
-  
-  *  **Organise live video to Parkes during demo. Tasso to talk to Shaun (Tasso) **  
- 
-  *  **Organise help for Python developments for demo displays (Chris/​Cormac)** 
- 
-NEXT OBSERVING SESSION 
- 
-  *  Some of the new 750 GB disks appear to have bigbuf problems at 512 Mbps. Claire to send Jim some disks for an out-of-session test at Hobart (Claire/​Jim). 11/2/09 - tests continue with high priority. Need to determine asap. 
- 
-  *  ** Observing support schedule for February session on the Wiki (Tasso)** 
- 
-  *  **Wiki entries for next session needed. To be constructed from schedules (Cormac)** 
-  *  **Remind PIS re late schedules. (Cormac) ** 
- 
-  *  Expedite ARCS support for data transfers (Steven/​Cormac/​Tasso/​Chris) 
- 
-  *  **Decision to ship disks from Nov 08 session to Curtin. (Parkes/​Narrabri)** 
-  *  **Observatories will need disk IDs to ship. (Claire)** 
- 
- 
-OTHER 
- 
-  *  **Need updated inventory of all XRAID disks. Letters to observatories and correlators. (Claire)** 
- 
-  *  **Mopra shows some data glitches from remote recording. Need to assess problem during correlation. (Hayley/​Claire/​Chris) ** 
- 
-  *  **Need determination of data format for tests with GMRT. Tasso to write to Adam Deller for information and support (Tasso) ** 
- 
- 
- 
- 
-  *  Next telecon: Wed 1 April, 2009, 2pm Sydney time (Tasso/​Chris) 
- 
  
lbaops/telecon2009feb11.txt · Last modified: 2015/12/18 16:39 (external edit)